# Exploring Lee-Yang and Fisher Zeros in the 2D Ising Model through Multi-Point Padé Approximants

### v2

@@ Abstract
We present a numerical calculation of the Lee-Yang and Fisher zeros of the 2D Ising model using multi-point Padé approximants. We perform simulations for the 2D Ising model with ferromagnetic couplings both in the absence and in the presence of a magnetic field using a cluster spin-flip algorithm. We show that it is possible to extract genuine signature of Lee Yang and Fisher zeros of the theory through the poles of magnetization and specific heat, using multi-point Padé method. We extract the poles of magnetization using Padé approximants and compare their scaling with known results. We verify the circle theorem associated to the well known behaviour of Lee Yang zeros. We present our finite volume scaling analysis of the zeros done at $T=T_{c}$ for a few lattice sizes, extracting to a very good precision the (combination of) critical exponents $\beta\delta$. The computation at the critical temperature is performed after the latter has been determined via the study of Fisher zeros, thus extracting both $\beta_{c}$ and the critical exponent $\nu$. Results already exist for extracting the critical exponents for the Ising model in 2 and 3 dimensions making use of Fisher and Lee Yang zeros. In this work, multi-point Padé is shown to be competitive with this respect and thus a powerful tool to study phase transitions.
@@ I Introduction
The knowledge of phase transitions is essential in determining the equation of state of a physical system. Phase transitions are characterised by non-analyticities that develop in thermodynamic functions in the limit of infinite degrees of freedom in a system. In finite volume systems, thermodynamic functions like the free energy are analytic functions of system parameters and the study of the onset of divergences becomes challenging. However, there are remnants of these divergences (if existing) even in finite volumes and occur as peaks in the susceptibility of the order parameter, whose height and width scale with volume. The conventional way to look for phase transitions in finite systems is to study the finite size scaling of these susceptibilities [1, 2, 3, 4, 5, 6]. Although these techniques are still used extensively in the study of phase transitions in the fields of condensed matter and lattice gauge theories, one of our goals in this paper is to show the power of an alternative scheme in extracting critical exponents for certain theories from their numerical simulations at finite volumes, namely, the Lee-Yang (LY) zero analysis [7, 8, 9, 10, 11]. While results on the numerical extraction of LY zeros from some lattice models already exist [12, 13, 14, 15, 16, 17, 18, 19], the novelty of our work is to extract them using only the leading order cumulants of the partition function, evaluated at multiple points in parameter space, using a rational function re-summation of the cumulants called the “multi-point Padé” method.
@@ II Complex zeroes of the partition function
where $\{a,b\ldots\}$ are configuration dependent coefficients mentioned above. The symmetry of the partition function under $z\to z^{-1}$ is the statement that the Hamiltonian in Eq. (1) is invariant under the combined action of reversing the external magnetic field and flipping all the spins.
@@ 3. III
    A phase transition is said to occur at some critical temperature $T_{crit}$ when these zeros approach the real axis of the external field parameter, in the thermodynamic limit. A first order phase transition is said to occur when these zeros cross the real axis. A second order transition is said to occur when the zeros, along with their complex conjugate pair pinch the real axis, in the thermodynamic limit.
@@
Therefore, for a finite $N$, the partition function is strictly positive when $\beta$ and $H$ are real. Hence, the only zeros of $Z_{N}$ occur in the complex plane of $z$ or on the $z<0$ axis. In the complex $H$ plane this translates to having only complex zeros of $Z_{N}(z(H))$. However, in the infinite volume limit, $Z_{N}$ is an infinite series which can now have non-trivial zeros on the $z>0$ axis, or real zeros in $H$. Since real zeros of the partition function mark the onset of phase transitions, we have recovered the well-known result that phase transitions cannot occur in a finite volume. The natural question to ask now will be on how to access these complex zeros of the partition function. This is fortunately not hard to answer because thermal cumulants, like the average magnetization in the case of the Ising model, are related to the derivative of the logarithm of the partition function (see Section V). This means that the zeros of $Z_{N}(H)$ will appear as poles of the average magnetization as a function of $H$ and the specific heat capacity as a function of $\beta$. The aim of the next few sections is to study in some detail these poles in the complex plane of the external magnetic field and inverse temperature. Arguments similar to those presented in this section can also be found in more detail in the existing literature [11], and we refer the reader to [15] for a detailed review on the subject.
@@ III.1 Simulation : To extract Fisher zeros
$C_{H}$ $=\left(\frac{\partial E}{\partial T}\right)_{H}=\beta^{2}\left(\langle E^{2}\rangle-\langle E\rangle^{2}\right)$ (4)
@@ III.2 Simulation : To extract Lee-Yang zeros
$\chi_{H}$ $=\left(\frac{\partial M}{\partial H}\right)_{T}=\beta\left(\langle M^{2}\rangle-\langle M\rangle^{2}\right)$ (5)
@@
The resulting $\langle M\rangle$ and $\chi_{H}$, per lattice site, are shown in Fig. 2 and the details of the number of configurations are as follows : For each lattice size, $625K$ configurations were used to estimate the average magnetization and the resulting susceptibility  [^4] Data generated from the simulations are available and will be made public via a separate data publication. Until then, it will be made available on request. We will now proceed to describe the multi-point Padé method for extracting the zeros of the partition function from the poles of the $C_{H}$ and $\langle M\rangle$ data.
Figure 2: Top : Average magnetization per site as a function of $H$ from simulations. Bottom : Susceptibility per site calculated from the configurations generated
@@ IV.1 The multi-point Padé method in a nutshell
In what follows, we will solve this system of linear equations to determine the coefficients of the polynomials $P_{m}$ and $Q_{n}$. Other techniques like a generalized $\chi^{2}$ method can also be used to estimate the coefficients of the rational functions by minimizing the distance between the measured Taylor coefficients and the required rational function, weighted by the errors available on the measured coefficients. This has been compared to the linear solver method in [20].
@@ IV.2 The method at work for the 2D Ising model
We will now focus on the results of the approximation and the singularity structure obtained from the Padé procedure outlined in the previous sub-section, for the average magnetization. We will first show the results of the approximation in Fig. 3. That we can see the rational functions approximate the data correctly is not surprising because we have essentially done a rational interpolation through the data, since our input only consisted of the average magnetization values and not the susceptibilities. Therefore, the first real success of the rational approximation is to see how well its derivatives approximate the susceptibilities, as shown in Fig. 4 for the $L=10\,\,,15$ data. This is a nice result because the rational function was constructed assuming only the knowledge of the zeroth order Taylor coefficients in the expansion of the average magnetization and it faithfully returns the expected first order coefficients at all the input points. Having gained some confidence in our multi-point Padé approximation, we can now proceed to look at the singularity structure, i.e. we will now study the zeros and poles of the rational function constructed in the complex $H$ plane  [^5] All results shown below have been obtained using the linear algebra routines of MATLAB [61] and all plots have been made using Jupyter notebooks [62].
@@
Selecting two lattice sizes at $T_{c}$, we refer the reader to Fig. 5. In the figure, we depict the zeros of the numerator (black pentagons) and of the denominator (red crosses) of $R^{m}_{n}(H)$ at different values of the lattice size $L$, i.e. $L=15$ (left panel) and $L=30$ (right panel). The pale blue points are shown to indicate the interval of points where the value of average magnetization was used as an input to build the rational function. We can easily make a couple of key observations.
@@
  • A few zeros of the denominator are canceled by corresponding zeros of the numerator. These are not genuine pieces of information: actually their location can vary when varying e.g. the order of the Padé approximant $[m,n]$. On the other hand, genuine pieces of information (i.e. actual zeros and poles) stay stable to a very good precision.
@@
  • In order to analyse the genuine poles more carefully, we remove the spurious zeros and poles from the singularity structure. We further remove all remaining zeros to focus the attention on poles and refer to Fig. 6. Looking at the pole structure of the figures, we can make the following comments : (1) As the lattice size $L$ gets larger, the closest singularity $H_{0}$ gets closer to the real axis, with real parts being consistent with $\mbox{Re}(H_{0})=0$. This means that all poles of the magnetization are purely imaginary in the $H$ variable. This is known in literature as the Circle theorem, because in the fugacity like variable, $z=e^{-\beta H}$, they lie on the unit circle. (2) Referring the reader to Section II, all poles occur with their complex conjugate pairs. (3) Although this observation is highly dependent on the statistics used, which for the purposes of this work is relatively high, we can see an increase in the number of zeros as the lattice sites are increases. All of these points seem to indicate that we are observing genuine Lee-Yang zeros.
@@ IV.3 On the stability of the closest poles to the real $H$ axis
  • Statistical errors: These are the errors propagated from the noise on the measured Taylor coefficients to the poles of the rational function approximant. The procedure used to estimate this error was to solve the system of linear equations in Eq. (7) repeatedly by choosing new Taylor coefficients for each solve. These coefficients are drawn from a Gaussian distribution centered at the central values of the measured coefficients and having standard deviation given by the noise on the corresponding Taylor coefficient. We refer the reader to Fig. 7, where the cloud of green points are the closest poles extracted for the $L=10$ lattice. The scatter is from $\sim 700$ noise-varied Taylor coefficients used to construct the respective rational functions. The error bars obtained from this analysis are used in the study of scaling in the next section.
  • Systematic errors: These are the errors on the closest poles resulting from varying the order of the multi-point Padé and (or) changing the selection of the input points used to construct the rational approximant, while keeping the order the same. We can estimate this error on the closest pole by repeatedly solving for various orders of the rational function, while working with the central values of the Taylor coefficients. The idea is to change the input points by deleting data in a systematic way to construct the rational function of varying orders. We now refer to Fig. 8 where we show the singularity structure for $L=15$ lattice  [^6] The reader may notice this is a slightly different plot than that of Fig. 5 (Top), although they are both using the same input data for the $L=15$ lattice (all available Taylor coefficients at their central values). The reason is that the rational functions have both been calculated using different methods : the one in Fig. 5 using the Linear solver and the one in this plot using the least square fitting procedure. The goal is to further show the stability of the genuine structure of the poles obtained. Here the scatter of poles arising from changing the order of the rational function is shown as dark blue points, notice that the scatter is over only around $\sim 50$ points, as the goal is only to show the stability of the zero.
Note that the systematic effects explained above only arise because of the noise present on the data. If the data had been clean, any order of the rational approximation would give identical results. In other words, the systematic errors mentioned above are correlated with the statistical errors.
All in all, from the error analysis performed above, the stability of the closest poles of average magnetization give us confidence in the fact that we are extracting genuine poles of the function, i.e. genuine zeros of the partition function and thus LY zeros. We can now proceed to analyse the scaling of these poles with the lattice volume to extract physical information like the critical exponents.
@@
Figure 4: Derivative of the rational function obtained in Fig. 3 for $L=10\,\,,15$ plotted against the susceptibility data. Note that this is not an interpolation and no data on the susceptibility was used in the construction of the rational function.
Figure 5: Zeros of the numerator (black pentagons) and of the denominator (red crosses) of the rational approximant $R^{m}_{n}(H)$ for the magnetisation on $L=15$ (Top) and $L=30$ (Bottom). The pale blue circles are the points used as input for the Padé. Notice that the closest singularity to the real axis gets closer to the real $H$ axis as $L$ gets larger, with real parts being consistent with $\mbox{Re}(H_{0})=0$.
Figure 6: Genuine poles extracted from the rational approximant of the average magnetization, shown for two lattice sizes (Top : $L=15$, Bottom : $L=30$). We want to highlight that these poles follow the circle theorem. They lie on the imaginary $H$ axis and the closest pole moves closer to the real $H$ axis as the lattice size increases. Moreover, the number of poles increases.
Figure 7: Stability of closest pole for $L=10$. The green cloud represents closest poles extracted from varying the Taylor coefficients with noise drawn from a Gaussian distribution. The singularity structure shown, red crosses and black pentagons, are for the central values of the data at $L=10$. This is the result of drawing the coefficients $\sim 700$ times

### v3

# Exploring Lee-Yang and Fisher Zeros in the 2D Ising model through multipoint Padé approximants

@@ Abstract
We present a numerical calculation of the Lee-Yang and Fisher zeros of the 2D Ising model using multi-point Padé approximants. We perform simulations for the 2D Ising model with ferromagnetic couplings both in the absence and in the presence of a magnetic field using a cluster spin-flip algorithm. We show that it is possible to extract genuine signature of Lee Yang and Fisher zeros of the theory through the poles of magnetization and specific heat, using multi-point Padé method. We extract the poles of magnetization using Padé approximants and compare their scaling with known results. We verify the circle theorem associated to the well known behaviour of Lee Yang zeros. We present our finite volume scaling analysis of the zeros done at $T=T_{c}$ for a few lattice sizes, extracting to a good precision the (combination of) critical exponents $\beta\delta$. The computation at the critical temperature is performed after the latter has been determined via the study of Fisher zeros, thus extracting both $\beta_{c}$ and the critical exponent $\nu$. Results already exist for extracting the critical exponents for the Ising model in 2 and 3 dimensions making use of Fisher and Lee Yang zeros. In this work, multi-point Padé is shown to be competitive with this respect and thus a powerful tool to study phase transitions.
@@ I Introduction
The knowledge of phase transitions is essential in determining the equation of state of a physical system. Phase transitions are characterised by non-analyticities that develop in thermodynamic functions in the limit of infinite degrees of freedom in a system. In finite volume systems, thermodynamic functions like the free energy are analytic functions of system parameters and the study of the onset of divergences becomes challenging. However, there are remnants of these divergences (if existing) even in finite volumes and occur as peaks in the susceptibility of the order parameter, whose height and width scale with volume. The conventional way to look for phase transitions in finite systems is to study the finite size scaling of these susceptibilities [1, 2, 3, 4, 5, 6]. Although these techniques are still used extensively in the study of phase transitions in the fields of condensed matter and lattice gauge theories, one of our goals in this paper is to show the power of an alternative scheme in extracting critical exponents for certain theories from their numerical simulations at finite volumes, namely, the Lee-Yang (LY) zero analysis [7, 8, 9, 10, 11]. While results on the numerical extraction of LY zeros from some lattice models already exist [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], the novelty of our work is to extract them using only the leading order cumulants of the partition function, evaluated at multiple points in parameter space, using a rational function re-summation of the cumulants called the “multi-point Padé” method.
@@ II Complex zeroes of the partition function
The symmetry of the partition function under $z\to z^{-1}$ is the statement that the Hamiltonian in Eq. (1) is invariant under the combined action of reversing the external magnetic field and flipping all the spins.
@@ 3. III
    A phase transition is said to occur at some critical temperature $T_{crit}$ when these zeros approach the real axis of the external field parameter, in the thermodynamic limit. The behaviour of the density of (Lee-Yang) zeros at the real axis is used to distinguish between a first and a second order phase transition. At a second order transition the complex conjugate pair gets infinitesimally close to the real axis but the density of zeros is zero on the real axis. On the other hand, a first order transition sees a non-zero density of zeros on the real axis.
@@
Therefore, for a finite $N$, the partition function is strictly positive when $\beta$ and $H$ are real. Hence, the only zeros of $Z_{N}$ occur in the complex plane of $z$ or on the $z<0$ axis. In the complex $H$ plane this translates to having only complex zeros of $Z_{N}(z(H))$. However, in the infinite volume limit, $Z_{N}$ is an infinite series which can now have non-trivial zeros on the $z>0$ axis, or real zeros in $H$. Since real zeros of the partition function mark the onset of phase transitions, we have recovered the well-known result that phase transitions cannot occur in a finite volume. The natural question to ask now will be on how to access these complex zeros of the partition function. This is fortunately not hard to answer because thermal cumulants, like the average magnetization in the case of the Ising model, are related to the derivative of the logarithm of the partition function (see Section V). This means that the zeros of $Z_{N}(H)$ will appear as poles of the average magnetization as a function of $H$ and the specific heat capacity as a function of $\beta$. The aim of the next few sections is to study in some detail these poles in the complex plane of the external magnetic field and inverse temperature. Arguments similar to those presented in this section can also be found in more detail in the existing literature [11, 30], and we refer the reader to [20] for a detailed review on the subject.
@@ III.1 Simulation : To extract Fisher zeros
$C_{H}$ $=\left(\frac{\partial\langle E\rangle}{\partial T}\right)_{H}=\beta^{2}\left(\langle E^{2}\rangle-\langle E\rangle^{2}\right)$ (4)
@@ III.2 Simulation : To extract Lee-Yang zeros
$\chi_{H}$ $=\left(\frac{\partial\langle M\rangle}{\partial H}\right)_{T}=\beta\left(\langle M^{2}\rangle-\langle M\rangle^{2}\right)$ (5)
@@
The resulting $\langle M\rangle$ and $\chi_{H}$, per lattice site, are shown in Fig. 2 and the details of the number of configurations are as follows: For each lattice size, $625K$ configurations were used to estimate the average magnetization and the resulting susceptibility. We will now proceed to describe the multi-point Padé method for extracting the zeros of the partition function from the poles of the $C_{H}$ and $\langle M\rangle$ data.
Refer to captionRefer to caption
Figure 2: Top : Average magnetization per site as a function of $H$ from simulations. Bottom : Susceptibility per site calculated from the configurations generated.
@@ IV.1 The multi-point Padé method in a nutshell
In what follows, we will solve this system of linear equations to determine the coefficients of the polynomials $P_{m}$ and $Q_{n}$ while restricting our analysis to diagonal ($[q,q]$) and near diagonal ($[q,q+1]$) type Padé approximants. Other techniques like a generalized $\chi^{2}$ method can also be used to estimate the coefficients of the rational functions by minimizing the distance between the measured Taylor coefficients and the required rational function, weighted by the estimated errors on the measured coefficients. This has been compared to the linear solver method in [25].
@@ IV.2 The method at work for the 2D Ising model
We will now focus on the results of the approximation and the singularity structure obtained from the Padé procedure outlined in the previous sub-section, for the average magnetization. We will first show the results of the approximation in Fig. 3. That we can see the rational functions approximate the data correctly is not surprising because we have essentially done a rational interpolation through the data, since our input only consisted of the average magnetization values and not the susceptibilities. Therefore, the first real success of the rational approximation is to see how well its derivatives approximate the susceptibilities, as shown in Fig. 4 for the $L=10\,\,,15$ data. This is a nice result because the rational function was constructed assuming only the knowledge of the zeroth order Taylor coefficients in the expansion of the average magnetization and it faithfully returns the expected first order coefficients at all the input points  [^4] Notice that in Fig. 4, around $H\sim 0.1$ we observe a small bump in the rational approximation of the susceptibility. This is due to a near cancellation of a zero-pole pair on the real $H$ axis (see Fig. 5, top). The reason these zero-pole pairs don’t cancel exactly sometimes is because the data comes with noise. For a more detailed explanation see [25].. Having gained some confidence in our multi-point Padé approximation, we can now proceed to look at the singularity structure, i.e. we will now study the zeros and poles of the rational function constructed in the complex $H$ plane  [^5] All results shown below have been obtained using the linear algebra routines of MATLAB [70] and all plots have been made using Jupyter notebooks [71].
@@
Selecting two lattice sizes at $T_{c}$, we refer the reader to Fig. 5. In the figure, we depict the zeros of the numerator (black pentagons) and of the denominator (red crosses) of $R^{m}_{n}(H)$ at different values of the lattice size $L$, i.e. $L=15$ (left panel) and $L=30$ (right panel). The order the rational approximant constructed is $[m,n]=[25,25]$. The pale blue points are shown to indicate the interval of points where the value of average magnetization was used as an input to build the rational function. We can easily make a couple of key observations.
@@
  • A few zeros of the denominator are canceled by corresponding zeros of the numerator. These are not genuine pieces of information: actually their location can vary when varying e.g. the order of the Padé approximant $[m,n]$. On the other hand, genuine pieces of information (i.e. actual zeros and poles) stay stable to a very good precision. Notice that the alternating sequence of zeros and poles along the imaginary $H$ axis is how a Padé approximant alludes to a branch cut eventually showing up in the thermodynamic limit  [^6] In principle we find other zeros and poles very far away from the region of interpolation, however these usually occur when we demand higher order of the Padé approximant and are not stable feature of the approximated function. This kind of structure even occurs when we try to approximate analytic functions..
@@
  • In order to analyse the genuine poles more carefully, we remove the spurious poles from the singularity structure. We further remove all remaining zeros to focus the attention on poles and refer to Fig. 6. Looking at the pole structure of the figures, we can make the following comments : (1) As the lattice size $L$ gets larger, the closest singularity $H_{0}$ gets closer to the real $H$ axis. We claim that we verify the circle theorem, which translates to the Lee-Yang zeros lying on the imaginary $H$ axis. However, looking at Fig. 5 the reader may notice that some of the poles for the $L=30$ lattice are shifted away from imaginary $H$ axis, and hence seemingly violating the $H\to-H$ symmetry. This is because we have used the central values of magnetization which were simulated at each quoted value of $H$ and hence do not respect the anti-symmetry exactly, but do so within errors. In support of our claim we provide the location of the genuine poles of magnetization at the central values along with error bars in Fig. 6, obtained by performing a bootstrap procedure, i.e. for a given point we extract new data from a Gaussian distribution with mean given by central value and standard deviation given by the estimated error. Notice that only the closest pole tends to appear stable. We further observe that for the smaller $L=15$ lattice, the second pole has some uncertainty easy to inspect by eye. For the $L=30$, we observe a third pole which now appears with even larger uncertainty. In the following, only the closest pole is important for the discussions of scaling. (2) Referring the reader to Section II, all poles occur with their complex conjugate pairs. (3) Although this observation is highly dependent on the statistics used, which for the purposes of this work is relatively high, we can see an increase in the number of zeros as the number of lattice sites increases. All of these points seem to indicate that we are observing genuine Lee-Yang zeros.
@@ IV.3 On the stability of the closest poles to the real $H$ axis
  • Statistical errors: These are the errors propagated from the estimated error on the measured Taylor coefficients to the poles of the rational function approximant. The procedure used to estimate this error was to solve the system of linear equations in Eq. (7) repeatedly by choosing new Taylor coefficients for each solve. These coefficients are drawn from a Gaussian distribution centered at the central values of the measured coefficients and having standard deviation given by the estimated error on the corresponding Taylor coefficient. We refer the reader to Fig. 7, where the cloud of green points are the closest poles extracted for the $L=10$ lattice. The scatter is from repeating the bootstrap procedure around $\sim 700$ times, keeping the order of the Padé approximant fixed at $[m,n]=[25,25]$.
  • Systematic errors: These are the errors on the closest poles resulting from varying the order of the multi-point Padé and (or) changing the selection of the input points used to construct the Padé approximant, using only the central values of the input data. The idea is to change the input points by deleting data in a systematic way to construct the rational function of varying orders. We vary the order of the Padé approximant from $[m,n]=[25,25]\to[5,5]$. We now refer to Fig. 8 where we show the singularity structure for $L=15$ lattice  [^7] The reader may notice this is a slightly different plot than that of Fig. 5 (Top), although they are both using the same input data for the $L=15$ lattice (all available Taylor coefficients at their central values). The reason is that the rational functions have both been calculated using different methods : the one in Fig. 5 using the Linear solver and the one in this plot using the least square fitting procedure. The goal is to further show the stability of the genuine structure of the poles obtained. Here the scatter of poles arising from changing the order of the rational function is shown as dark blue points, notice that the scatter is over only around $\sim 50$ points, as the goal is only to show the stability of the closest pole.
Note that the systematic errors mentioned above are correlated with the statistical errors. All in all, from the error analysis performed above, the stability of the closest poles of average magnetization gives us confidence in the fact that we are extracting genuine poles of the function, i.e. genuine zeros of the partition function and thus LY zeros. We can now proceed to analyse the scaling of these poles with the lattice volume to extract physical information like the critical exponents.
Refer to captionRefer to caption
@@
Figure 4: Derivative of the rational function obtained in Fig. 3 for $L=10\,\,,15$ plotted against the susceptibility data. Note that this is not an interpolation and no data on the susceptibility was used in the construction of the rational function. Since this is a derivative of an $[m,n]=[25,25]$ rational function, the order is $[m,n]=[24,25]$
Refer to captionRefer to caption
Figure 5: Zeros of the numerator (black pentagons) and of the denominator (red crosses) of the rational approximant $R^{m}_{n}(H)$ for the magnetisation on $L=15$ (Top) and $L=30$ (Bottom), with $[m,n]=[25,25]$. The pale blue circles are the points used as input for the Padé. Notice that the closest singularity to the real axis gets closer to the real $H$ axis as $L$ gets larger, with real parts being consistent with $\mbox{Re}(H_{0})=0$.
Refer to captionRefer to caption
Figure 6: Genuine poles extracted from Fig. 5 along with their error bars, shown for two lattice sizes (Top : $L=15$, Bottom : $L=30$). We want to highlight that these poles follow the circle theorem. They lie on the imaginary $H$ axis (within errors) and the closest pole moves closer to the real $H$ axis as the lattice size increases. Moreover, the number of poles increases.
Figure 7: Stability of closest pole for $L=10$. The green cloud represents closest poles extracted from varying the Taylor coefficients with noise drawn from a Gaussian distribution. The singularity structure shown, red crosses and black pentagons, are for the central values of the data at $L=10$. This is the result of drawing the coefficients $\sim 700$ times using a rational function of order $[m,n]=[25,25]$